OpenStudy (anonymous):

For the following function, f, determine the domain, locate the critical points of f; use the First Derivative Test to locate local extrema; identify any absolute extrema; determine the intervals for which f is increasing/decreasing; determine the intervals for which f is concave up/down; identify and inflection points; determine any aspymptotes; and sketch the function. f(x) = x ------ x^2 - 9

OpenStudy (anonymous):

\[f(x) = \frac{ x }{ x ^{2} - 9 }\]

OpenStudy (tkhunny):

What's your plan? The Domain is an eyeball problem. You should just look at it and shout it out. ALL REAL NUMBERS except x = 3 and x = -3. Tell me where I got that so easily.

OpenStudy (anonymous):

yah, I got that - you just set the denominator equal to zero. I have the derivative but need help with critical points. 3 and -3 are not critical points and when I set the numerator of the derivative to zero i get an imaginary number.

OpenStudy (tkhunny):

Okay, since Critical Points for a rational function are REAL numbers, what does that say about the critical points as exposed by the 1st Derivative?

OpenStudy (anonymous):

that they dont exist? ... this is where im confused

OpenStudy (tkhunny):

There is nowhere on f(x), for REAL x, where the 1st derivative is zero. Perfect. Let's move on to the 2nd Derivative.

OpenStudy (anonymous):

The second derivative is used to find the inflection points and use them to find where concavity changes.

OpenStudy (tkhunny):

Excellent. Let's see it.

OpenStudy (anonymous):

okay, so \[f \prime(x) = \frac{ -x ^{2} - 9 }{ (x ^{2} -9)^{2} }\] then \[f \prime \prime(x) = \frac{ (x ^{2}-9)(-2x) - (-x ^{2} -9)(4x(x ^{2}-9)) }{ (x ^{2} -9)^{4}}\] right?

OpenStudy (tkhunny):

So close! Make that first \((x^{2} - 9)\) squared.

OpenStudy (anonymous):

oops! It is on my paper.. . just didnt type it that way :)

OpenStudy (anonymous):

would inflection points be -3,3,0?

OpenStudy (tkhunny):

Very good. I get \(f"(x) \ \dfrac{2x(x^{2}+ 27)}{(x^{2} - 9)^{3}}\) Tell me why x = 3 and x = -3 can be inflection points. They are not in the Domain!

OpenStudy (anonymous):

because they are points where change if it is decreasing or increasing even though it may not be differentiable

OpenStudy (tkhunny):

Very good answer. If it's positive on one side and negative on the other, something must have happened.

OpenStudy (anonymous):

and then I just need to plug in numbers to the second derivative on the intervals (below) to find of they are positive or negative, which will give me the concavity? \[(-\infty, -3)(-3, 0) (0, 3) (3, \infty) \]

OpenStudy (tkhunny):

So we have for Concavity \(Downward < -3 < Upward < 0 < Downward < 3 < Upward\) Okay, we already have the x- and y-intercepts (0,0). We have only an horizontal asymptote and we're ready to draw.

OpenStudy (anonymous):

perfect :) Thank you so much! This all seemed so easy when it was separate and easy example problems in class and in my notes!

OpenStudy (tkhunny):

One piece at a time. Just wade through it systematically. Very, very good work.

OpenStudy (anonymous):

thanks again :)